OpenStudy (anonymous):

how is cot inverse x = to tan inverse 1/x

OpenStudy (anonymous):

anyone got something?

hartnn (hartnn):

ok do you know tan x = 1/cot x ?

OpenStudy (anonymous):

yeah

hartnn (hartnn):

cot inv x = tan inv 1/x means we need to prove tan (cot inv x) = 1/x

hartnn (hartnn):

tan (cot inv x) = 1/ cot(cot inv x) and what is cot (cot inv x) = ... ?

OpenStudy (anonymous):

1/x? lol

hartnn (hartnn):

cot (cot inv x) is just x! :)

OpenStudy (anonymous):

this is the first trig stuff we come across all year still getting warmed up ahah

hartnn (hartnn):

okk. so did u get how cot inverse x = to tan inverse 1/x ?

OpenStudy (anonymous):

question what did you do to set tan(cot inv x)=1/cot(cot inv x)

hartnn (hartnn):

tan theta = 1/ cot theta

hartnn (hartnn):

theta = cot inv x

hartnn (hartnn):

more doubts?

OpenStudy (anonymous):

no that makes sense...

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^